OpenStudy (mtalhahassan2):

Graphing each parabola, labelling the vertex and axis of symmetry. y = -3(x+4) square 2 +2 Is anyone tell me how can we do this question with the help of table of value

OpenStudy (mtalhahassan2):

hey

OpenStudy (anonymous):

you cant do that

OpenStudy (aum):

Is the equation \(y = -3(x+4)^2 + 2\) ?

OpenStudy (anonymous):

no

OpenStudy (mtalhahassan2):

yeah this equation y=−3(x+4)2+2 is right

OpenStudy (aum):

Okay. Do you have to do this by creating a table of x and y values? There is a easier way to graph this equation than create a table.

OpenStudy (mtalhahassan2):

ok then tell me the easy way

OpenStudy (aum):

\(y = x^2\) is the equation of a standard parabola, with vertex at the origin and the parabola opens upward. \(y = -x^2\) is is the equation of a standard parabola, with vertex at the origin and the parabola opens downward.